# CAST: Contrastive Adaptation and Distillation for Semi-Supervised Instance Segmentation

### v3

@@ Abstract
Instance segmentation demands costly per-pixel annotations and large models. We introduce CAST, a semi-supervised knowledge distillation (SSKD) framework that compresses pretrained vision foundation models (VFM) into compact experts using limited labeled and abundant unlabeled data. CAST unfolds in three stages: (1) domain adaptation of the VFM teacher(s) via self-training with contrastive pixel calibration, (2) distillation into a compact student via a unified multi-objective loss that couples standard supervision and pseudo-labels with our instance-aware pixel-wise contrastive term, and (3) fine-tuning on labeled data to remove residual pseudo-label bias. Central to CAST is an instance-aware pixel-wise contrastive loss that fuses mask and class scores to mine informative negatives and enforce clear inter-instance margins. By maintaining this contrastive signal across both adaptation and distillation, we align teacher and student embeddings and fully leverage unlabeled images. On Cityscapes and ADE20K, our 11$\approx$ 11$\times$ 11 × smaller student surpasses its adapted VFM teacher(s) by +3.4 AP (33.9 vs. 30.5) and +1.5 AP (16.7 vs. 15.2) and outperforms state-of-the-art semi-supervised approaches.
@@ Appendix 1 Introduction
Pixel-level instance segmentation is notoriously expensive: annotating detailed masks can take hours per image, and training state-of-the-art detectors often requires hundreds of GPU hours, putting many applications out of reach [12, 15]. At the same time, recent advancements in VFMs [22, 20, 43, 19] have significantly elevated the capabilities of computer vision systems, enabling exceptional performance across diverse perception benchmarks [2].
Motivation. Despite remarkable achievements, foundation models still cannot serve specific downstream tasks sufficiently well due to two major issues: (1) the heavy computational overhead during deployment making these models impractical for environments with limited resources [37]; and (2) their inherently generic nature, which leads to suboptimal performance on tasks that demand domain specific expertise. The latter stems from foundation models being optimized to perform well across a wide variety of tasks, rather than being finely tuned for the nuanced requirements of specialized applications [5]. This challenge is particularly prominent in applications involving outdoor environments, such as autonomous driving, and indoor settings, such as robotic perception [13]. SSKD for instance segmentation seeks to compress large models into efficient student models by leveraging both limited labeled data and abundant unlabeled images. Current distillation methods either treat VFMs as fixed feature extractors with simple pseudo-labeling or focus on coarse semantic tasks, failing to exploit the rich structure of large unlabeled datasets to refine per-pixel predictions. Consequently, adjacent instances remain poorly separated and accuracy degrades sharply under scarce labels. We address these issues by adapting VFMs via self-training to enhance pseudo-label fidelity, and by injecting an instance-aware pixel-wise contrastive loss that leverages unlabeled data to enforce clear inter-instance margins, yielding sharper masks and superior performance in the low-label regime.
Status quo. Knowledge distillation has evolved from task-agnostic compression [16, 8] to adapting VFMs for downstream tasks. For classification and semantic segmentation, Vemulapalli et al. [30] distill a VFM matching its output on an unlabeled transfer set, and SAM-CLIP [31] fuses CLIP and SAM. However, neither method targets per-pixel instance masks nor exploits dense self-supervision from the unlabeled pool. Pure semi-supervised instance segmentation methods, such as [17, 3] train teachers from scratch, doubling GPU cost, and still produce blurred boundaries under scarce labels. To our knowledge, no prior work unifies VFM adaptation, unlabeled data-driven pixel-wise refinement, and extreme student compression for high fidelity instance segmentation.
@@
  1. 1.
    We introduce an instance-aware pixel-wise contrastive loss that fuses mask and class predictions to drive stronger inter-instance separation, and show how to sample negatives efficiently in a instance centric setting.
  2. 2.
    We propose CAST, a SSKD pipeline with three phases: (i) self-training a foundation teacher with contrastive calibration, (ii) distilling into a compact student using supervised, pseudo label, and pixel-wise contrastive objectives, and (iii) final supervised finetuning to remove bias thereby unifying supervised, semi-supervised, and self-supervised signals.
  3. 3.
    We conduct extensive experiments on Cityscapes and ADE20K, demonstrating that our 11$\approx$ 11$\times$ 11 × smaller student surpass self-trained teacher by +3.4 AP (33.9 vs. 30.5) and +2.1 AP (16.7 vs. 14.6), and outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
@@ Appendix 2 Related Work
Vision Foundation Models. Vision foundation models (VFMs)  [22, 20, 25, 41, 4] have revolutionized computer vision through large scale pretraining. In parallel, recent trends focus on combining VFMs to extend their capabilities [26, 43]. While these models excel in open-set recognition and transfer learning, their computational demands yet hinder edge deployment. Recent efforts merge VFMs via distillation: Wang et al [31] unify SAM and CLIP via multi-task learning, while Zhang etal [44] distill CLIP and DINOv2 into a compact model with moderate scale data distillation. We extend these paradigms by leveraging VFMs for instance segmentation, focusing on balancing robustness with computational efficiency.
@@
Knowledge Distillation in Vision. Knowledge distillation (KD) has become a ubiquitous technique to transfer knowledge from teachers with high capacity to lightweight students for efficient deployment. Early methods distilled softened logits or intermediate features [16] in a task-agnostic way, while later feature-based approaches capture structured spatial cues (e.g., pixel-wise similarity, channel distributions) [23, 27]. Modern methods tackle VFMs’ scale and complexity: [28, 38] distills VFMs to impart zero-shot and multimodal capabilities, further multi teacher approaches [18, 39] combine complementary expertise. Vemulapalli et al. [30] adapt a VFM to the target task and then distill on a large unlabeled set for classification and semantic segmentation. Overall with these advances in vision KD We build our method on the insight that a strong teacher (or ensemble of teachers) can guide a lightweight instance segmentation model to high performance. our approach explicitly integrates semi-supervised learning and pixel-level contrastive losses for instance segmentation, to focus on bridging the gap between rich representation of VFMs and compact, efficient student networks.
@@
Semi-Supervised Learning. Self-training (or pseudo-labeling) has become a foundational paradigm in semi-supervised learning (SSL), where a model leverages its own predictions with high confidence and iteratively refine itself [35]. This approach has proven effective across vision tasks, improving image classification performance [35] and boosting object detection accuracy when annotation budgets are tight [21]. To counteract error accumulation from noisy pseudo-labels [29] use exponential moving average of label predictions, or [6] employ curriculum labeling schemes that gradually incorporate harder examples. More recent work applies pseudo-labeling for large pretrained models through targeted finetuning and adaptive pseudo selection strategies [14]. While many SSL methods focus on classification or detection, several have extended this method to dense prediction tasks [10, 40].
@@
We study self-training with self-supervised contrastive learning and task-specific adaptation. Global contrastive frameworks such as SimCLR [7], MoCo [9], and their detection extensions [34] established the value of large scale visual discrimination learning. Further per pixel contrastive approaches [32, 36, 46, 33, 1] have shown promise in retaining spatial sensitivity though they yet conflate pixels from different instances of the same class. We extend these advances by synergizing self-training and self-supervised contrastive learning, and introduce a novel instance aware negative sampling strategy designed specifically for the demands of instance segmentation.
@@
Figure 1: CAST framework overview. Top: Three-stage pipeline: (1) adapt a pretrained VFM teacher to the target domain via self-training with contrastive pixel calibration; (2) transfer knowledge to a compact student using instance-aware pixel-wise contrastive sampling; (3) fine-tune the student on labeled data to correct residual pseudo-label bias. Bottom: Detailed view of stage (2): fused mask and class score maps produce anchor pixels, which are sampled across weak/strong views to form positive and negative pairs; an MLP projects features and computes the pixel-wise contrastive loss. Dashed arrows denote no gradient flow; red components are trainable, blue are frozen.
@@ 3.1 Overview
In semi-supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
@@
3.2 Pixel-wise Contrastive Loss
@@
Let $z^{\rm weak},z^{\rm strong}\in\ R^{B\times N\times D}$ be $\ell_{2}$-normalized embeddings from two views of each image, where $B$ is the number of images in one mini batch, $N=h\times w$ the number of pixels, and $D$ the embedding dimension. For each pixel $p\in{1,2,...,N}$ and image index $b\in{1,...,B}$, the corresponding embedding vector is denoted as $z_{b,p}\in\mathbb{R}^{D}$. We construct the positive pair by sampling the weak and strong embeddings for each pixel. The positive similarity between the two views is
@@
The pixel-wise contrastive loss is then the standard NT-Xent over all anchors:
@@
Debiased Pixel-Level Negative Sampling.
To mine true inter-instance pairs without quadratic cost, we derive a per pixel sampling distribution by fusing mask and class probabilities. Let $M\in\mathbb{R}^{B\times K\times H\times W}$, and $L\in\mathbb{R}^{B\times K\times(C+1)},$ be the model’s mask and class logits respectively. We first resize $M$ to the feature resolution $(h\times w)$ and then normalize logits to probability distributions $P_{m}$ and $P_{c}$ via softmax along instance and class dimensions respectively.
@@
We score any two pixels $(b,p)\neq(b^{\prime},q)$ by $\tilde{y}$ being $\ell_{2}$-normalized vector of pseudo probability map.
[…]
@@
Theoretical Insight. To give a formal rationale for augmenting our pixel-wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
@@ 3.3 CAST Framework
Teacher adaptation. Starting from pretrained weights $\theta_{T}^{0}$, we first fine-tune on the labeled set $\mathcal{D}^{l}$:
[…]
We then generate pseudo-labels $\hat{y}_{j}^{u}=f_{\theta_{T}^{\prime}}(x_{j}^{u})$, reset to $\theta_{T}^{0}$and fine-tune on $\mathcal{D}^{l}\cup\{\!(x_{j}^{u},\hat{y}_{j}^{u})\!\}$:
[…]
This two-step contrastive calibration yields a specialized teacher whose pseudo-labels are both accurate and spatially consistent for the target domain.
@@
[…]
This pass removes pseudo-label drift and sharpens decision boundaries for in-domain data.
@@ 4.1 Experimental Protocol
For our semi-supervised experiments, we randomly sample $10\%$ of the Cityscapes training set as labeled and use the remaining images as unlabeled. For ADE20K, to reduce computational cost we use stratified sampling to select $20\%$ of the training pool while preserving the original per category distribution, yielding 3,537 images. From this subset, 1,000 images (10 per instance class) serve as labeled data and the remaining 2,537 as unlabeled. Detailed statistics are provided in Appendix A, and we will release the exact split files alongside our code.
Teacher and Student Architectures. Our teacher is a fused ensemble of Grounding-DINO-Large [20] and SAM2-L [25]. Since the official Grounding-DINO is closed-source, we use its open-source counterpart mm-Grounding-DINO [45]. For the student, we pair a DINOv2-S encoder [22] with a DPT-S decoder head [24], followed by a lightweight transformer decoder module in the spirit of Mask2Former [11]. Our choice of the DINOv2+DPT backbone is motivated by the recent successes of “Depth AnythingV2” in monocular depth estimation [41] and UniMatchV2 in semantic segmentation [42], and aims to facilitate future multimodal fusion work. We evaluate the impact of different student designs in Sec. 4, and defer the complete optimizer, learning rate schedules, and other hyperparameters to Appendix B.
Implementation Details. All experiments were conducted on Ubuntu 22.04 with Python 3.10 and PyTorch 2.6.0 (CUDA 12.6). Teacher adaptation runs were executed on 2× NVIDIA A100 GPUs, while student training and fine-tuning runs used 2× NVIDIA GeForce RTX 4090 GPUs. For instance, a single fine-tuning run of the teacher (Grounding-DINO) on the supervised Cityscapes split required $\approx 3.5$ GPU hours; a single student training run for this dataset took $\approx 17$ GPU hours without the pixel-wise contrastive branch and $\approx 22$ GPU hours with it. All reported results in the following tables are averages over three independent runs with different random seeds.
@@ 4.2 Main Results
We extensively evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state-of-the-art semi-supervised baselines, and benchmark them against our CAST pipeline. Table 1 reports maskAP and AP50 on Cityscapes and ADE20K. In the teacher adaptation stage (568 M parameters), adding our pixel-level contrastive loss boosts Cityscapes maskAP from 29.8 to 30.5 (+0.7) and AP50 from 54.9 to 56.6 (+1.7); on ADE20K, maskAP rises from 14.6 to 15.2 (+0.6) and AP50 from 23.6 to 24.5 (+0.9). These improvements confirm that pixel-wise supervision sharply improves feature discrimination and reduces pseudo-label noise. Figure 2 plots key pipeline efficiency metric on a logarithmic scale for both teacher and student models.
Figure 2: Comparison of Teacher and Student Model efficiency Metrics (Log Scale).
In the student distillation stage (52 M parameters, just 9 % of the teacher), our semi-supervised model with pixel-level loss achieves 32.2 maskAP and 56.5 AP50 on Cityscapes, outperforming the Guided dist. baseline by +1.4 maskAP. After fine-tuning, the student reaches 33.9 maskAP (+3.4 over the best teacher) and 58.7 AP50. On ADE20K, it records 16.1 maskAP and 27.4 AP50 semi-supervised, and 16.7 maskAP (+1.5) with 28.0 AP50 after fine-tuning, underscoring CAST’s robustness across domains.
Table 1: Instance segmentation results on Cityscapes and ADE20K. We report teacher-adaptation (568 M params) and student distillation (52 M params) under each method.
Method Data Regime Cityscapes ADE20K
maskAP AP50 maskAP AP50
Teacher Adaptation
[…]
Supervised Labeled only $28.7$ $53.4$ $14.2$ $23.5$
Self-training [35] Labeled + Unlabeled $29.7$ $54.9$ $14.6$ $23.6$
+ EMA-Augmented [21] Labeled + Unlabeled $29.8$ $54.9$ $14.8$ $23.7$
+ pixel-level loss (ours) Labeled + Unlabeled 30.5 56.6 15.2 24.5
Student Distillation
[…]
PAIS [17] Labeled + Unlabeled $22.9$ $44.9$ $10.3$ $18.3$
Guided dist. [3] Labeled + Unlabeled $30.8$ $52.9$ $14.2$ $23.8$
+ unlabeled KD [30] Unlabeled only $24.4$ $45.6$ $5.1$ $9.3$
+ labeled+unlabeled KD (ours) Labeled + Unlabeled $30.7$ $54.9$ $14.4$ $25.2$
+ pixel-level loss (ours) Labeled + Unlabeled 32.2 56.5 16.1 27.4
Fine-tuned (best) Labeled only 33.9 58.7 16.7 28.0

### v4

@@ Abstract
Instance segmentation demands costly per-pixel annotations and computationally expensive models. We introduce CAST, a semi-supervised knowledge distillation (SSKD) framework that compresses pre-trained vision foundation models (VFM) into compact experts using limited labeled and abundant unlabeled data. CAST unfolds in three stages: (1) domain adaptation of the VFM(s) via self-training with contrastive calibration, (2) knowledge transfer through a unified multi-objective loss, and (3) student refinement to mitigate residual pseudo-label bias. Central to CAST is an instance-aware pixel-wise contrastive loss that fuses mask and class scores to extract informative negatives and enforce clear inter-instance margins. By maintaining this contrastive signal across both adaptation and distillation, we align teacher and student embeddings and fully leverage unlabeled images. On Cityscapes and ADE20K, our $\approx 11\times$ smaller student improves over its zero-shot VFM teacher(s) by +8.5 and +7.1 AP, surpasses adapted teacher(s) by +3.4 and +1.5 AP, and further outperforms state-of-the-art SSKD methods on both benchmarks.
@@ Appendix 1 Introduction
Pixel-level instance segmentation is notoriously expensive: annotating detailed masks can take hours per image, and training state-of-the-art detectors often requires hundreds of GPU hours, putting many applications out of reach [12, 15]. Recent advancements in vision foundation models (VFMs) [22, 20, 45, 19] have substantially expanded the capabilities of computer vision systems, achieving strong performance across diverse perception benchmarks [2].
Motivation. Despite remarkable achievements, foundation models still cannot serve specific downstream tasks sufficiently well due to two major issues: (1) the heavy computational overhead during deployment making these models impractical for environments with limited resources [38]; and (2) their inherently generic nature, which leads to suboptimal performance on tasks that demand domain specific expertise [28]. The latter stems from foundation models being optimized to perform well across a wide variety of tasks, rather than being finely tuned for the nuanced requirements of specialized applications [5]. This challenge is prominent in applications that involve outdoor environments, such as autonomous driving, and indoor settings, such as robotic perception [13]. Semi-supervised knowledge distillation (SSKD) for instance segmentation seeks to compress large models into efficient student models by leveraging both limited labeled data and abundant unlabeled images. Current distillation methods either treat VFMs as fixed feature extractors with simple pseudo-labeling or focus on coarse semantic tasks, failing to exploit the rich structure of unlabeled datasets to refine per-pixel predictions. Consequently, adjacent instances remain poorly separated and accuracy degrades sharply under scarce labels. We address these issues by adapting VFMs via self‐training to enhance pseudo-label fidelity, and by injecting an instance-aware pixel‐wise contrastive loss that leverages unlabeled data to enforce clear inter‐instance margins, yielding sharper masks and superior performance in the low‐label regime.
Status quo. Knowledge distillation has evolved from task‐agnostic compression [16, 8] to adapting VFMs for downstream tasks. For classification and semantic segmentation, Vemulapalli et al. [31] distill a VFM matching its output on an unlabeled transfer set, and SAM-CLIP [32] fuses CLIP and SAM. However, neither method targets per-pixel instance masks nor exploits dense self‐supervision from the unlabeled pool. Pure semi‐supervised instance segmentation methods, such as [17, 3] train teachers from scratch, doubling GPU cost, and still produce noisy masks under scarce labels. To our knowledge, no prior work unifies VFM adaptation, unlabeled data‐driven pixel‐wise refinement, and extreme student compression for instance segmentation.
@@
  • We introduce an instance-aware pixel-wise contrastive loss that fuses mask and class predictions to drive stronger inter-instance separation, and show how to sample negatives efficiently in an instance centric setting.
  • We propose CAST, a SSKD pipeline with three phases: (i) adapting the foundation teacher via self-training with contrastive calibration, (ii) distilling into a compact student using a unified objective that combines supervised, pseudo-label, and pixel-wise contrastive losses, and (iii) supervised fine-tuning to reduce residual bias, unifying supervised, semi-supervised, and self-supervised signals.
  • We conduct extensive experiments on Cityscapes and ADE20K, demonstrating that our $\approx 11\times$ smaller student improves over its zero-shot VFM teacher(s) by +8.5 and +7.1 AP, surpasses adapted teacher(s) by +3.4 and +1.5 AP, and further outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
@@ Appendix 2 Related Work
Vision Foundation Models. VFMs [22, 20, 25, 42, 4] have revolutionized computer vision through large scale pre-training. In parallel, recent trends focus on combining VFMs to extend their capabilities [26, 45]. While these models excel in open-set recognition and transfer learning, their computational demands yet hinder edge deployment. Recent efforts merge VFMs via distillation: Wang et al. [32] unify SAM and CLIP via multi-task learning, while Zhang et al. [46] distill CLIP and DINOv2 into a compact model with data distillation. We extend these paradigms by leveraging VFMs for instance segmentation, focusing on balancing robustness with computational efficiency.
@@
Knowledge Distillation in Vision. Knowledge distillation (KD) has become a ubiquitous technique to transfer knowledge from teachers with high capacity to lightweight students for efficient deployment. Early methods distilled softened logits or intermediate features [16] in a task-agnostic way, while later feature-based approaches capture structured spatial cues (e.g., pixel-wise similarity, channel distributions) [23, 27]. Modern methods tackle VFMs’ scale and complexity: [29, 39] distills VFMs to impart zero-shot and multimodal capabilities, further multi-teacher approaches [18, 40] combine complementary expertise. Vemulapalli et al. [31] adapt a VFM to the target task and then distill on a large unlabeled set for classification and semantic segmentation. Building on these advances in vision knowledge distillation, we posit that a strong teacher (or ensemble of teachers) can effectively guide a lightweight instance segmentation model to high performance. Our approach explicitly integrates semi-supervised learning and pixel-level contrastive signals for instance segmentation, to focus on bridging the gap between rich representation of VFMs and compact, efficient student networks.
@@
Semi-Supervised Learning. Self‐training (or pseudo‐labeling) has become a foundational paradigm in semi‐supervised learning (SSL), where a model leverages its own predictions with high confidence and iteratively refines itself [36]. This approach has proven effective across vision tasks, improving image classification performance [36] and boosting object detection accuracy when annotation budgets are tight [21]. To counteract error accumulation from noisy pseudo‐labels [30] use exponential moving average of label predictions, or [6] employ curriculum labeling schemes that gradually incorporate harder examples. More recent work applies pseudo-labeling for large pre-trained models through targeted finetuning and adaptive pseudo selection strategies [14]. While many SSL methods focus on classification or detection, several have extended this method to dense prediction tasks [10, 41].
@@
We study self‐training with self‐supervised contrastive learning and task‐specific adaptation. Global contrastive frameworks such as SimCLR [7], MoCo [9], and their detection extensions [35] established the value of large-scale visual discrimination learning. Further per-pixel contrastive approaches [33, 37, 48, 34, 1] have shown promise in retaining spatial sensitivity though they yet conflate pixels from different instances of the same class. We extend these advances by synergizing self-training and self-supervised contrastive learning, and introduce a novel instance-aware negative sampling strategy designed specifically for the demands of instance segmentation.
@@
Figure 1: CAST framework overview. Top: Three-stage pipeline: (1) adapt a pre-trained VFM teacher to the target domain via self-training with pixel-level contrastive calibration; (2) distill knowledge into a compact student using instance-aware contrastive sampling; (3) fine-tune the student on labeled data to correct residual pseudo-label bias. Bottom: Detailed view of stage (2): fused mask and class score maps produce anchor pixels, sampled across weak/strong views to form positive/negative pairs; an MLP projects features for the contrastive loss. Dashed arrows denote no gradient flow; red modules are trainable, blue are frozen.
@@ 3.1 Overview
In semi‐supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
@@
3.2 Pixel‐wise Contrastive Loss
@@
Let $z^{\rm weak},z^{\rm strong}\in\ R^{B\times N\times D}$ be $\ell_{2}$‐normalized embeddings from two views of each image, where $B$ is the number of images in one mini batch, $N=h\times w$ the number of pixels, and $D$ the embedding dimension. For each pixel $p\in{1,2,...,N}$ and image index $b\in{1,...,B}$, the corresponding embedding vector is denoted as $z_{b,p}\in\mathbb{R}^{D}$. We construct the positive pair by sampling the weak and strong embeddings for each pixel. The positive similarity between the two views is
@@
The pixel‐wise contrastive loss is then the standard NT‐Xent over all anchors:
@@
Debiased Pixel‐Level Negative Sampling.
To mine true inter‐instance pairs without quadratic cost, we derive a per pixel sampling distribution by fusing mask and class probabilities. Let $M\in\mathbb{R}^{B\times K\times H\times W}$, and $L\in\mathbb{R}^{B\times K\times(C+1)},$ be the model’s mask and class logits respectively. We first resize $M$ to the feature resolution $(h\times w)$ and then normalize logits to probability distributions $P_{m}$ and $P_{c}$ via softmax along instance and class dimensions respectively.
@@
We score any two pixels $(b,p)\neq(b^{\prime},q)$ by $\tilde{y}$ being $\ell_{2}$‐normalized vector of pseudo probability map.
$s^{\mathrm{deb}}\bigl((b,p),(b^{\prime},q)\bigr)=\max\bigl(0,\;1-\langle\tilde{y}[b,p],\,\tilde{y}[b^{\prime},q]\rangle\bigr),$
@@
Theoretical Insight. To give a formal rationale for augmenting our pixel‐wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
@@ 3.3 CAST Framework
Teacher adaptation. Starting from pretrained weights $\theta_{T}^{0}$, we first fine‐tune on the labeled set $\mathcal{D}^{l}$:
$\theta_{T}^{\prime}=\arg\min_{\theta}\;\mathcal{J}\bigl(\theta;\,\mathcal{D}^{l},\varnothing;\,0,\lambda_{\rm pxl}\bigr).$
We then generate pseudo‐labels $\hat{y}_{j}^{u}=f_{\theta_{T}^{\prime}}(x_{j}^{u})$, reset to $\theta_{T}^{0}$and fine-tune on $\mathcal{D}^{l}\cup\{\!(x_{j}^{u},\hat{y}_{j}^{u})\!\}$:
$\theta_{T}^{\prime\prime}=\arg\min_{\theta}\;\mathcal{J}\bigl(\theta;\,\mathcal{D}^{l},\mathcal{D}^{u};\,1,\lambda_{\rm pxl}\bigr).$
This two‐step contrastive calibration yields a specialized teacher whose pseudo-labels are both accurate and spatially consistent for the target domain.
@@
$\theta_{s}^{\dagger}=\arg\min_{\theta_{s}^{*}}\;\mathcal{J}\bigl(\theta_{s}^{*};\,\mathcal{D}^{l},\varnothing;\,0,0\bigr),$
This pass removes pseudo-label drift and sharpens decision boundaries for in‐domain data.
@@ 4.1 Experimental Protocol
[…]
Teacher and Student Architectures. Our teacher is a fused ensemble of Grounding-DINO-Large [20] and SAM2-L [25]. Since the SOTA model of Grounding-DINO is closed-source, we use its open-source counterpart mm-Grounding-DINO [47]. For the student, we pair a DINOv2-S encoder [22] with a DPT-S decoder head [24], followed by a lightweight transformer decoder module in the spirit of Mask2Former [11]. Our choice of the DINOv2+DPT backbone is motivated by the recent successes of “Depth AnythingV2” in monocular depth estimation [42] and UniMatchV2 in semantic segmentation [43], and aims to facilitate future multimodal fusion work. We evaluate the impact of different student designs in Sec. 4, and defer the complete optimizer, learning rate schedules, and other hyperparameters to Appendix B.
@@ 4.2 Main Results
We evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state‐of‐the‐art semi‐supervised baselines, and benchmark them against our CAST pipeline. Table 1 reports maskAP and maskAP50 on Cityscapes and ADE20K. In the teacher adaptation stage (568M parameters), adding our pixel‐level contrastive loss boosts Cityscapes maskAP from 29.7 to 30.5 (+0.8) and maskAP50 from 54.9 to 56.6 (+1.7); on ADE20K, maskAP rises from 14.6 to 15.2 (+0.6) and maskAP50 from 23.6 to 24.5 (+0.9). These improvements confirm that pixel‐wise supervision sharply improves feature discrimination and reduces pseudo-label noise.
In the student distillation stage, our 52M-parameter student (9% of the composite teacher model) achieves 32.2 maskAP and 56.5 maskAP50 on Cityscapes with pixel-level loss, outperforming prior SOTA SSKD models. After fine-tuning, the student reaches 33.9 maskAP (+3.4 over the best teacher) and 58.7 maskAP50. On ADE20K, it attains 16.1 maskAP and 27.4 maskAP50 in the semi-supervised setting, and improves further to 16.7 maskAP (+1.5) and 28.0 maskAP50 after fine-tuning, underscoring CAST’s robustness across benchmarks. Additional ablations under varied label splits are presented in Section 4. To compare efficiency, Figure 2 plots key pipeline efficiency metric on a logarithmic scale for both teacher and student models.
Table 1: Main results on Cityscapes and ADE20K with 10% labeled data. We report teacher adaptation (568M) and student distillation (52M). * denotes adapted methods. Rows in gray are ours.
Method Data Regime Cityscapes ADE20K maskAP maskAP50 maskAP maskAP50 Teacher Adaptation Zero-shot VFM None (pretrained) $22.0$ $42.3$ $8.1$ $18.2$ Supervised fine-tuning Labeled only $28.7$ $53.4$ $14.2$ $23.5$ Self-training* [36] Labeled+Unlabeled $29.7$ 54.9 $14.6$ $23.6$ Unbiased Teacher* [21] Labeled+Unlabeled 29.8 54.9 14.8 23.7 CAST (teacher adaptation) Labeled+Unlabeled 30.5 56.6 15.2 24.5 Student Distillation Supervised fine-tuning Labeled only $21.1$ $38.7$ $13.9$ $24.2$ PAIS [17] Labeled+Unlabeled $22.9$ $44.9$ $10.3$ $18.3$ Guided dist. [3] Labeled+Unlabeled $30.8$ $52.9$ $14.2$ $23.8$ Vemulapalli et al.* [31] Unlabeled only $24.4$ $45.6$ $5.1$ $9.3$ CAST (knowledge transfer) Labeled+Unlabeled 32.2 56.5 16.1 27.4 CAST (student refinement) Labeled only 33.9 58.7 16.7 28.0
Figure 2: Efficiency comparison (log scale).